. Given the following information:

Spot rate €.6/$, 6 month forward rate €.61/$, interest rates p.a. in the US (Germany) are 4% (6%) respectively, a call option with a strike price of €.59/$ with a 2% premium, a put option with a strike price €.605/$ with a premium of 2.5%. How could you hedge a €200 million receivable? What is the breakeven exchange rate between the option alternative you choose and the forward market hedge?
Answer #1

HEDGING € 200million EUROS RECEIVABLE

MONEY MARKET HEDGE:

6 months interest in Germany =(6/2)%=3%=0.03

STEP 1 :Borrow Present Value of € 200million =200/1.03= €   194.17 million in Germany

STEP2: Convert into US dollar at current spot rate of € 0.6/$

Amount received =194.17/0.6= $323.62million

STEP 3: Invest 323.62 million at six monthly interest of (4/2=)2%

Amount Received at end of six months =323.62*1.02= $330.10 million

STEP4: Amount of € 200millon received after 6 months will be used to clear the debt in EURO including interest

OPTION HEDGE

The Risk is increase in spot rate . With increasing spot rate dollar received will be less.

Hence CALL option to be used to hedge the risk

BUY 200 million EURO call option with strike price= € 0.59/$

Cost of Option =2%*200 million EURO=4 million EURO=(4/0.6=)=$6.6666667 million

If 6.666667 million is borrowed in US,

Future Value with interest =6.6666667*1.02=$6.8 million

Net amount received after 6 months=((200/0.59)million-6.8 million)US dollars

Net amount received after 6 months=$332.1831 million

Since the amount received is higher than Money Market Hedge,

OPTION HEDGING WOULD BE USED.

Break Even exchange rate in Forward market:

200/332.1831=0.6021 EURO/Dollar

€   0.6021 /$